OpenStudy (armyrangers):

The first challenge to complete is the Penny Problem. The radio station is giving the winner of this challenge a prize pack that includes tickets to see his or her favorite band in concert. To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? You must show all work to receive credit.

OpenStudy (armyrangers):

@Blu-Girl

OpenStudy (anonymous):

Volume of penny * no. of pennies = volume of jar no. of pennies = volume of jar / volume of penny n = \[(\pi R ^{2} H)/ (\pi r ^{2}h)\] n = \[(\ R ^{2} H)/ (\ r ^{2}h)\] R= 6 inches H = 11.5 inches r = 0.375 h = 0.061

OpenStudy (armyrangers):

Thanks so much

OpenStudy (anonymous):

your welcome